OpenStudy (anonymous):

HELP!! PLEASE HELP ME SOLVE THIS I'M STILL VERY CONFUSED!! One focus is at (-12.92,-8). Find the other focus for the ellipse defined by this equation: ((x+7)^2)/(36)+((y+8)^2)/(1)=1

OpenStudy (primeralph):

http://www.mathsisfun.com/geometry/ellipse.html

OpenStudy (anonymous):

That just confused me even more.

OpenStudy (sirm3d):

the center of the ellipse lies midway the two foci. identify the coordinates of the center, plot it, and the given focus.

OpenStudy (littlenugget):

Here is the equation of an ellipse: \[ \frac{ x ^{2} }{ a ^{2}} + \frac{ y ^{2} }{ b ^{2} } =1 \] In the equation, a is bigger than b so we know that the ellipse is horizontal (left/right) We're going to graph this to make it easier. ;) Your equation: \[\frac{ (x+7)^{2} }{ 36 } + \frac{ (y+8)^{2} }{ 1 } = 1\]

OpenStudy (anonymous):

Okie dokie! :)

OpenStudy (littlenugget):

http://www.mathopenref.com/coordgeneralellipse.html At this website, jump down to where it says Ellipses not centered at the origin So for you equation, the 'h' is -7, and the 'k' is -8. And the center for an ellipse is (h,k). so when you plug it in, it's (-7,-8)  So now we have the center =) next, we need to find the axes! Yeah! So we know that 'a' is always the bigger one. In the equation, we have a^2. so we need to solve for that a. a^2=36 square root it a=6 now you know that the a is 6, you will next to go left and right 6 from the center. So let's solve for b now: b^2=1 square root it b=1 now you know that the b is 1, you will go up and down 1 from the center. I'll graph this. (sorry i'm taking so long)

OpenStudy (littlenugget):

 Here what the ellipse looks like, now i'm going to plug in the focus they give us: (-12.92,-8)